# GenAI vs. Human Fact-Checkers: Accurate Ratings, Flawed Rationales

### v1

@@ Abstract.
Despite recent advances in understanding the capabilities and limits of generative artificial intelligence (GenAI) models, we are just beginning to understand their capacity to assess and reason about the veracity of content. We evaluate multiple GenAI models across tasks that involve the rating of, and perceived reasoning about, the credibility of information. The information in our experiments comes from content that subnational U.S. politicians post to Facebook. We find that GPT-4o, one of the most used AI models in consumer applications, outperforms other models, but all models exhibit only moderate agreement with human coders. Importantly, even when GenAI models accurately identify low-credibility content, their reasoning relies heavily on linguistic features and “hard” criteria, such as the level of detail, source reliability and language formality, rather than an understanding of veracity. We also assess the effectiveness of summarized versus full content inputs, finding that summarized content holds promise for improving efficiency without sacrificing accuracy. While GenAI has the potential to support human fact-checkers in scaling misinformation detection, our results caution against relying solely on these models.
@@ 1. Introduction
Generative Artificial Intelligence (GenAI) has transformed workflows across industries and academia since its surge in popularity in 2022 (Abels, 2023). In scientific research, its adoption has accelerated at an unprecedented pace (Zhao et al., 2023), demonstrating transformative potential in diverse applications. Researchers have successfully employed GenAI models to impersonate survey respondents (Bisbee et al., 2024), generate counterfactual images for experimental research (Davidson, 2024), annotate text with human-comparable accuracy(Gilardi et al., 2023), identifying conspiracy theories (Diab et al., 2024), and preemptively debunk election misinformation (Linegar et al., 2024). Among these diverse applications of GenAI, a fundamental question remains about its ability to assess content credibility.
@@ 2. Method
Data: We constructed our dataset based on original data collected and presented by Biswas et al. (Biswas et al., 2024). They collect over 493k posts shared on Facebook by 5,147 state legislators between 2020 and 2021. They identify posts that share unreliable URLs. This dataset is particularly well-suited for our study, as the shared posts by these politicians are political texts, making them highly relevant to misinformation-related social harms.
@@ 3. Results
Performance Evaluation–Q1: The results from the hand-coded sample revealed that a majority (65%) of the low-credibility posts labeled by the domain were genuinely misleading or low-credibility. Performance was evaluated using macro F1 and Matthews Correlation Coefficient (MCC) metrics, averaged across all prompts for each zero-shot model. Results are presented in Table 1. GPT-4o demonstrated superior performance with a macro F1 of 0.82 and an MCC of 0.66 when rating combined content. MCC scores are generally lower because they provide a balanced evaluation across all components of the confusion matrix and are less sensitive to class imbalance (Chicco and Jurman, 2020). Following GPT-4o, Gemma2 achieved a macro F1 score of 0.80 and an MCC of 0.61. As an open source GenAI model with a relatively smaller parameter size, Gemma2-9b exhibits a comparable performance to the closed source-GPT 4o in rating content accuracy. This suggests that certain open-source GenAI models can be viable alternatives to proprietary models for tasks like content accuracy rating. Regardless of relative better performance of GPT-4o and Gemma2-9b, big variation in performance exists. For example, Flan-T5-XL struggled to assess content reliability with a macro F1 of 0.36 and an MCC of 0.12, which might be attributed to its high sensitivity to prompt design (Wei et al., 2021) or its architecture as an encoder-decoder (seq2seq) model designed for structured tasks.
@@
Can Summaries Help? –Q2: On summarized content, Gemma2-9b outperformed others with a macro F1 score of 0.69 and an MCC of 0.49, followed by LLAMA3-70B with a macro F1 score of 0.64 and an MCC of 0.44, as shown in Table 2. For the remaining models, F1 scores ranged from 0.34 for Flan-T5-XL to 0.57 for GPT-4o. As shown in Figure 1, the agreement scores with human coders are generally low, with the hightest score of 0.42 by Gemma2-9b and lowest score of 0.03 by Flan-T5X-XL. The small variation in both performance and agreement scores across models may be attributed to the limited length of the summarized content.
@@
As shown in Figure 2, for most models, an increase in the Rouge-1 threshold generally resulted in a decrease in mismatch rates. Beyond a threshold of 0.5, the mismatch rate stabilized. For instance, the mismatch rate for the Llama3.1-70B model decreased steadily from 66.67% at a threshold of 0.1 to 40.09% at a threshold of 1.0, indicating consistent improvement with higher thresholds. In contrast, Gemma2-9B exhibited a sharp rise in mismatch rates to 63.15% at a threshold of 0.2 before decreasing and stabilizing around 37.92% for thresholds of 0.6 and higher. At these lower thresholds, the model may be misclassifying a larger number of less confident or ambiguous summaries, leading to a higher mismatch rate. When summaries are of higher quality, the performance of models in detecting misinformation using summaries improves. Among all models, Gemma2-9B demonstrated superior summarization capacity, followed by Llama3.1-70B, making them strong candidates for assisting credibility ratings efficiently.
@@
The performance of three groups of reasoning show a similar pattern in Figure 5. The “hard” criteria by content quality present a balanced performance in the predictions of both reliable and low-credibility content, with slightly better performance for low-credibility predictions. This suggests that “hard” criteria based on verifiable aspects of content remain reliable. Language objectivity struggled with distinguishing reliable content. Emotive language and subjective framing further complicated accurate predictions of reliable content. The group of legislative behavior is less effective in detecting low-credibility content. Many public officials publicly share misleading or false information in their speeches.

### v3

@@ Abstract.
Despite recent advances in understanding the capabilities and limits of generative artificial intelligence (GenAI) models, we are just beginning to understand their capacity to assess and reason about the veracity of content. We evaluate multiple GenAI models across tasks that involve the rating of, and reasoning about, the credibility of information. The information in our experiments comes from content that subnational U.S. politicians post to Facebook. We find that GPT-4o, one of the most used AI models in consumer applications, outperforms other models, but all models exhibit only moderate agreement with human coders. Importantly, even when GenAI models accurately identify low-credibility content, their reasoning relies heavily on linguistic features and “hard” criteria, such as the level of detail, source reliability and language formality, rather than an understanding of veracity. We also assess the effectiveness of summarized versus full content inputs, finding that summarized content holds promise for improving efficiency without sacrificing accuracy. While GenAI has the potential to support human fact-checkers in scaling misinformation detection, our results caution against relying solely on these models.
@@ 1. Introduction
Generative Artificial Intelligence (GenAI) has transformed workflows across industries and academia since its surge in popularity in 2022 (Abels, 2023). In scientific research, its adoption has accelerated at an unprecedented pace (Zhao et al., 2023), demonstrating transformative potential in diverse applications. Researchers have successfully employed GenAI models to impersonate survey respondents (Bisbee et al., 2024), generate counterfactual images for experimental research (Davidson, 2024), annotate text with human-comparable accuracy(Gilardi et al., 2023), identify conspiracy theories (Diab et al., 2024), and preemptively debunk election misinformation (Linegar et al., 2024). Among these diverse applications of GenAI, a fundamental question remains about its ability to assess content credibility.
@@ 2. Method
Data: We constructed our dataset based on original data collected and presented by Biswas et al. (Biswas et al., 2024). They collected over 493k posts shared on Facebook by 5,147 state legislators between 2020 and 2021. To focus on the most misleading content, we refined the “unreliable” category (Tai et al., 2023), originally from Media Bias/Fact Check’s (MBFC) ratings, by selecting URL domains coded as low or very low factual to create a low-factual reference. Using this reference, we identified low-factual posts. The low-factual post dataset is particularly well-suited for our study, as the shared posts by these politicians are political texts, making them highly relevant to misinformation-related social harms.
@@ 3. Results
Performance Evaluation–Q1: The results from the hand-coded sample revealed that a majority (65%) of the low-credibility posts labeled by the domain were genuinely misleading or low-credibility. Performance was evaluated using macro F1 and Matthews Correlation Coefficient (MCC) metrics, averaged across all prompts for each zero-shot model. Results are presented in Table 1. GPT-4o demonstrated superior performance with a macro F1 of 0.82 and an MCC of 0.66 when rating combined content. MCC scores are generally lower because they provide a balanced evaluation across all components of the confusion matrix and are less sensitive to class imbalance (Chicco and Jurman, 2020). Following GPT-4o, Gemma2 achieved a macro F1 score of 0.80 and an MCC of 0.61. As an open source GenAI model with a relatively smaller parameter size, Gemma2-9b exhibits a comparable performance to the closed source-GPT 4o in rating content accuracy. This suggests that certain open-source GenAI models can be viable alternatives to proprietary models for tasks like content accuracy rating. Regardless of relative better performance of GPT-4o and Gemma2-9b, big variation in performance exists. For example, Flan-T5-XL struggled to assess content reliability with a macro F1 of 0.36 and an MCC of 0.12, which could be attributed to its high sensitivity to prompt design (Wei et al., 2021) or its architecture as an encoder-decoder (seq2seq) model designed for structured tasks.
@@
Can Summaries Help? –Q2: On summarized content, Gemma2-9b outperformed others with a macro F1 score of 0.69 and an MCC of 0.49, followed by LLAMA3-70B with a macro F1 score of 0.64 and an MCC of 0.44, as shown in Table 2. For the remaining models, F1 scores ranged from 0.34 for Flan-T5-XL to 0.57 for GPT-4o. As shown in Figure 1, the agreement scores with human coders are generally low, with the hightest score of 0.42 by Gemma2-9b and the lowest score of 0.03 by Flan-T5X-XL. The small variation in both performance and agreement scores across models may be attributed to the limited length of the summarized content.
@@
As shown in Figure 2, for most models, an increase in the Rouge-1 threshold generally resulted in a decrease in mismatch rates. Beyond a threshold of 0.5, the mismatch rate stabilized. For example, the mismatch rate for the Llama3.1-70B model decreased steadily from 66.67% at a threshold of 0.1 to 40.09% at a threshold of 1.0, indicating consistent improvement with higher thresholds. In contrast, Gemma2-9B exhibited a sharp rise in mismatch rates to 63.15% at a threshold of 0.2 before decreasing and stabilizing around 37.92% for thresholds of 0.6 and higher. At these lower thresholds, the model may be misclassifying a larger number of less confident or ambiguous summaries, leading to a higher mismatch rate. When summaries are of higher quality, the performance of models in detecting misinformation using summaries improves. Among all models, Gemma2-9B demonstrated superior summarization capacity, followed by Llama3.1-70B, making them strong candidates for assisting credibility ratings efficiently.
@@
The performance of three groups of reasoning show a similar pattern in Figure 5. The “hard” criteria by content quality present a balanced performance in the predictions of both reliable and low-credibility content, with slightly better performance for low-credibility predictions. This suggests that “hard” criteria based on verifiable aspects of content remain reliable. Language objectivity struggled with distinguishing reliable content. Emotive language and subjective framing further complicate accurate predictions of reliable content. The group of legislative behavior is less effective in detecting low-credibility content. Many public officials publicly share misleading or false information in their speeches.